OpenStudy (anonymous):

3. Explain how a four-term polynomial is factored by grouping and when a quadratic trinomial can be factored using this method. Include examples in your explanation

OpenStudy (jack1):

look, inkyvoyd has given you all of the "how too" and step by steps for solving up to here, but the below equation is a similar example, for reference, if it also helps? \[\large x^2+bx+20\] use FOIL to factorise \[\large (x+J)(x+K)\] so\[\large J × K=20\] and \[\large J + K=b\] so possible answers for J and K are 4 and 5, 2 and 10, 1 and 20 lets use 4 and 5, jus coz proof: \[\large (x+5)(x+4)\] now expand using FOIL First = \[\large x×x=x^2\] Outer = \[\large x×4=4x\] Inner = \[\large 5×x=5x\] Last = \[\large 5×4=20\] so full equation if we use this is \[\large x^2+4x+5x+20 →x^2+9x+20\] so b = 9x, and is made up of the terms 4x and 5x follow so far? now, if that equation is changed to : \[\large 2x^2+bx+20\] ... using FOIL, we still factorise to: \[\large (x+L)(2x+M)\] so\[\large L × M=20\] and \[\large 2L + M=b →(\text {as right bracket is 2x + M})\] so possible answers for L and M are 4 and 5, 2 and 10, 1 and 20 lets use 4 and 5 again, just coz proof: \[\large (x+5)(2x+4)\] now expand using FOIL First = \[\large x×2x=2x^2\] Outer = \[\large x×4=4x\] Inner = \[\large 5×2x=10x\] Last = \[\large 5×4=20\] so full equation if we use this is \[\large 2x^2+4x+10x+20 →2x^2+14x+20\] so b = 14x, and is made up of the terms 4x and 10x does this example help at all? @BasedGod1122 ???

OpenStudy (jack1):

alternatively, if b in \[x^2+bx+20\] = b in \[2x^2+bx+20\] then we sub in our original b value of 9, and solve 2nd equation \[2x^2+bx+20\] using quadratic formula